OpenStudy (anonymous):

So 2x +y =3 isn't a direct variation? Neither is 2.4+6=1.2y. But 8x+2y=0 is and y= 4x is the constant? And x-3y=0 is a direct variation and y=-1/3x is the constant.

OpenStudy (solomonzelman):

It is not a direct variation

OpenStudy (anonymous):

someone give me a medal!

OpenStudy (solomonzelman):

Direct variations look like y=5x or like y=-2x

OpenStudy (anonymous):

So it doesnt matter if you can graph it? You can graph any equation right? So a direct variation just has an x and y with something multiplied to x? So x-3y=0. Where y= -3x. And 8x+2y=0 where y=4x

OpenStudy (anonymous):

Those are but these are not----- 2x+y=3 isn't. Neither is 2.4x + 6 = 1.2y isn't either?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.